# A try for dark energy in quantum field theory: The vacuum energy of neutrino field

### v4

@@ Abstract
The quartic-divergent vacuum energy poses an ultraviolet (UV) challenge (the cosmological constant problem) in probing the nature of dark energy. Here we try to evaluate the contribution of the vacuum energy to dark energy with a method of the UV-free scheme. The result indicates that it is not a problem in the UV region but a question of the contributions of heavy fields being suppressed. Then, we propose an effective description via scale decoupling. The parameter spaces suggest that the active vacuum energy of neutrino fields can naturally meet the observation of dark energy density, and a neutrino with a typical mass $\sim$ 10 meV is expected. The normal ordering neutrinos are preferred by naturalness, and the neutrino mass window set by dark energy is 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV.
@@ I Introduction
Physics thrives on crisis Weinberg:1988cp , and the ultraviolet (UV) problem is one of the keys in the development of modern physics. In the context of the UV catastrophe in classical physics, the energy quanta were introduced by Planck Planck:1901tja while studying blackbody radiation problems, which led to the birth of quantum theory. In quantum field theory (QFT), the problem of UV divergences reappears, that is, loop corrections of a transition process are often UV divergent. It means that the transition amplitude obtained by Feynman rules (the physical input) is not directly equal to the physical result (the physical output). To extract the finite physical result from a UV-divergent input, the traditional approach is regularization & renormalization, that is, the UV divergence is firstly expressed by the regulator based on equivalent transformation and then removed by counterterms, with the mathematical structure of this route being $\infty$-$\infty$ - ∞. This traditional approach is successful in dealing with logarithmic divergences in the standard model (SM). When we leave the logarithmic region and move on, there are UV problems for power-law divergences, i.e.,
@@
(b) The non-renormalizable Einstein gravity. When one tries to quantize Einstein gravity, an infinite number of counterterms are required for graviton loops exhibiting power-law divergences.
@@
How do we deal with the above three UV problems? The traditional approach exhibits UV limitations in handling power-law divergences. In this case, the free flow of ideas is important. Dirac DiracTheEO viewed renormalization as a mathematical framework requiring refinement to address UV divergences; the physical laws chosen by nature were found to possess mathematical beauty. Feynman Feynman:1986er emphasized that mathematical self-consistency and rigor are essential in addressing UV divergences. In the vision of Dyson Dyson:1949ha and Schwinger Schwinger , future approaches to loop corrections would be based on finite quantities. Wilsonian effective theory Wilson:1971bg decouples scales by encoding UV effects into low-energy parameters via a physical cutoff, enabling systematic organization of loop contributions. Hence, if we go forward, we might have to consider introducing new physical ideas and new mathematical techniques to connect UV divergent inputs to the physical reality. For a new route, the requirement is that it first reproduces the successful renormalization results for logarithmic divergences, and then it can give an interpretation to the three UV problems caused by power-law divergences. For UV divergences of loops, assuming the physical contributions being local and insensitive to UV regions, a method of the UV-free scheme Jia:2023dub is introduced to derive loop corrections, which can reproduce the renormalization results for logarithmic divergences in SM. This method takes a route of analytic continuation from the physical input $\mathcal{T}_{\mathrm{F}}$ to the physical output $\mathcal{T}_{\mathrm{P}}$ with a mathematical structure $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$,
@@
Input{Equivalent transformation,Analytic $\mathrm{Input}\bigg{\{}\!\!\begin{array}[]{c}\textit{Equivalent transformation}\,,\infty-\infty\\ \textit{Analytic continuation}\,,\,\,\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}{}{}{}\end{array}\!\!\bigg{\}}\mathrm{Output}\,,$ (3)
@@
and UV divergences are systematically eliminated through $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$ without counterterms. Loop corrections in this method are constructed from local finite quantities that maintain scale decoupling, while ensuring UV-safe loop contributions through locality-based analytic continuation. To power-law divergences of loops, this method provides an interpretation to the hierarchy problem of Higgs mass within SM Jia:2023dub , and a try for its application to Einstein gravity (graviton loops) was investigated in Ref. Jia:2024lyy . Here another power-law divergence, the vacuum energy, is of our concern. The cosmological constant (dark energy) problem is an open fundamental challenge in physics today Weinberg:1988cp ; SupernovaSearchTeam:1998fmf ; SupernovaCosmologyProject:1998vns ; Peebles:2002gy ; Copeland:2006wr . In exploring the nature of dark energy, an unavoidable question is the contribution of the vacuum energy density $\rho_{\mathrm{vac}}$ that corresponds to a cosmological constant ($\Lambda=8\pi G\rho_{\mathrm{vac}}$). Since the form of the vacuum energy of a field ($\varpropto\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\sqrt{k^{2}+m^{2}}$) is similar to the integral of a loop, we will try it in the UV-free scheme in the following.
@@ II The vacuum energy in QFT
where $j_{i}$ is the spin of the field, and $g_{i}$ is the number of degrees of freedom of the field. This integral is a quartic divergence. Here the UV-free scheme Jia:2023dub ; Jia:2024lyy is adopted to evaluate the physical contribution of Eq. (4), which is a physical analytic continuation of the integral containing UV divergences. The Feynman amplitude $\mathcal{T}_{\mathrm{F}}^{i}=\rho_{0}^{i}$ is the physical input, and the Feynman-like amplitude $\mathcal{T}_{\mathrm{F}}^{i}(\xi)$ here can be obtained with $\sqrt{k^{2}+m_{i}^{2}}$ replaced by $\sqrt{k^{2}\!+m_{i}^{2}\!+\xi}$. The physical transition amplitude $\mathcal{T}_{\mathrm{P}}^{i}$, i.e., the physical output, is
@@
In the limit $m_{i}\to 0$, the result is zero for a massless field. [^1] For another type UV-divergence caused by the discrete sequence, i.e., the Casimir effect, the Riemann zeta function $\zeta(-1)=-\frac{1}{12}$ can be adopted as the analytic continuation. The energy scale $\mu_{\Lambda}$ is considered as a scale describing the character of vacuum energy, like the strong interaction scale $\Lambda_{\mathrm{QCD}}$ and the electroweak scale $v$.
Here is an interpretation of scales in QFT. The physical result of total tree-level and loop-level contributions in a scattering process is independent of the choice of an energy scale $\mu$, and the total physical result can be equivalently described by a tree-level process with an effective coupling at an effective scale $\mu_{\mathrm{eff}}$. The energy scale $\mu_{\Lambda}$ of the vacuum energy is similar to the scale $\mu_{\mathrm{eff}}$ in the total physical result, which is different from the free parameter $\mu$ in order-by-order expansions. In addition, $\mu_{\mathrm{eff}}$ changes with the scattering energy, while the scale $\mu_{\Lambda}$ describes the reference scale for the ground states of all fields, i.e., all zero-point fields have the same reference scale $\mu_{\Lambda}$. Moreover, $\mu_{\Lambda}$ may be a fixed or slow-running scale with the expansion of the universe.
@@ III The ripple description and neutrino field
Let us consider the contribution of a field to dark energy. The result of Eq. (6) indicates that it is no longer a problem in the UV region. The actual question is revealed, that is, the contributions of heavy fields should be suppressed by some unknown mechanisms that are beyond the current understanding of QFT. If we analyze the question with Wilsonian framework Wilson:1971bg , to vacuum energy, there may be scale decoupling between microscopic quantity and macroscopic effect. We will try to give it an effective description in the view of scale decoupling, that is, the total zero-point energy is the microscopic energy foundation as a background, and quantum fluctuations with cumulative contributions characterized by $\mu_{\Lambda}$ manifest macroscopic effects through fluctuation-carrying fields. The scale $\mu_{\Lambda}$ is a watershed scale to a carrier field with a mass $m_{i}$, that is, the contribution of a field’s vacuum energy to dark energy is active, inert in the case of $m_{i}\lesssim\mu_{\Lambda}$, $m_{i}\gg\mu_{\Lambda}$, respectively. This picture likes ripples on a lake, necessitating $\mu_{\Lambda}-$characterized fluctuation factor for ground states of fields (e.g., the energy quanta distribution in blackbody radiation research). This is the ripple description, as depicted in Fig. 1. Considering the Hubble parameter $H_{0}\approx$ 70 km s-1 Mpc-1 and dark energy contributing 71% of the critical density $3H_{0}^{2}/8\pi G$, the dark energy density is about $(2.3~{}\mathrm{meV})^{4}$ (1 meV $=10^{-3}$ eV), and $\mu_{\Lambda}$ is roughly estimated to be $10-100$ meV. Thus, the dark energy density is sensitive to the neutrino field, especially the neutrino mass. [^2] Here is more discussions about possible neutrino contributions, e.g., the relic neutrinos with varying masses Fardon:2003eh ; Fardon:2005wc ; Takahashi:2005kw , the slow-roll quintessence with extended neutrino sector Hung:2000yg ; Barbieri:2005gj , the cosmological quantum vacuum energy of axion and neutrino deVega:2023fvw , etc. Considering Eq. (6) and dark energy density, a typical neutrino mass is $m_{\nu}^{\mathrm{tp}}\sim 2.3\times(32\pi^{2})^{\frac{1}{4}}\approx$ 9.7 meV (here $|\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}|^{\frac{1}{4}}\sim 1$). As a rough prediction, to produce the required dark energy density, there should be a neutrino with a mass $\sim$ 10 meV.
@@
Now we turn to neutrino masses. The masses of three neutrinos are $m_{1}$, $m_{2}$ and $m_{3}$, and the mass differences between them can be extracted by neutrino oscillations. The mass-splitting results ParticleDataGroup:2024cfk are $\Delta m_{21}^{2}\approx 7.41\times 10^{-5}\mathrm{eV}^{2}$, $\Delta m_{31}^{2}\sim\Delta m_{32}^{2}\approx 2.437\times 10^{-3}\mathrm{eV}^{2}$ (normal ordering) and $\Delta m_{31}^{2}\sim\Delta m_{32}^{2}\approx-2.498\times 10^{-3}\mathrm{eV}^{2}$ (inverted ordering). Considering the naturalness with an effective mass $m_{\nu}^{\mathrm{tp}}$, the quasidegenerate spectrum with $m_{1}\simeq m_{2}\simeq m_{3}\gg\sqrt{|\Delta m_{32}^{2}|}$ is not favored. It is possible to give an estimate on the mass spectrum of neutrinos. Here we consider a specific implementation of the ripple description by a test fluctuation factor $e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ in Gaussian distribution to characterize the active contribution of a field (another possible form $e^{-m_{i}/\mu_{\Lambda}}$ in Boltzmann distribution is not preferred by the naturalness of parameter spaces), and $g_{i}$ together with this factor can be taken as an effective active degrees of freedom $g_{i}^{\ast}$ with $g_{i}^{\ast}=g_{i}e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ ($\rho_{\mathrm{vac}}=\sum_{i}(-1)^{2j_{i}}g_{i}^{\ast}\frac{m_{i}^{4}}{64\pi^{2}}\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}$). For a field with a mass $m_{i}\gg\mu_{\Lambda}$, its contribution to the dark energy density is negligible due to the tiny value of $g_{i}^{\ast}$ (this is similar to the freezing of effective degrees of freedom of heavy particles at a low temperature in the early universe). The hierarchical spectra of neutrinos for normal ordering and inverted ordering with the required dark energy density are shown in Fig. 2 and Fig. 3, respectively. In Fig. 3, considering the naturalness of $\mu_{\Lambda}$, the inverted ordering is not favored. For the normal ordering shown in Fig. 2, there is a region that neutrino masses at given $\mu_{\Lambda}$ can naturally produce the active vacuum energy required by the dark energy density, i.e., the parameter spaces forming a mass-energy window. Likewise, the neutrino mass window set by dark energy is 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV, and the total neutrino mass is 67.5 meV $\lesssim m_{1}+m_{2}+m_{3}\lesssim$ 87.4 meV. This mass window can be tested by the future experiments.
@@
Let us give a brief discussion about the Hubble tension. A slightly higher $H_{0}=(73.04\pm 1.04$) km s-1 Mpc-1 Riess:2016jrr ; Riess:2021jrx from the nearby universe and a slightly lower $H_{0}=(67.4\pm 0.5$) km s-1 Mpc-1 Planck:2018vyg from the early universe indicate a possible Hubble tension (the tension may be alleviated Freedman:2024eph ). For $H_{0}\approx$ 67.4 km s-1 Mpc-1 and $H_{0}\approx$ 73 km s-1 Mpc-1, the corresponding dark energy densities are about $(2.24~{}\mathrm{meV})^{4}$ and $(2.37~{}\mathrm{meV})^{4}$, respectively. If the energy scale $\mu_{\Lambda}$ has a slow-running behavior and traverses the upper part of the mass-energy window in Fig. 2, the Hubble tension can be explained by this slow-running behavior, i.e., a running cosmological constant. If $\mu_{\Lambda}$ is fixed, new explanations and mechanisms need to be taken into account.

### v5

@@ Abstract
The quartic-divergent vacuum energy poses an ultraviolet (UV) challenge (the cosmological constant problem) in probing the nature of dark energy. Here we try to evaluate the contribution of the vacuum energy to dark energy with a method of the UV-free scheme. The result indicates that it is not a problem in the UV region but a question of the contributions of heavy fields being suppressed. Then, we explore an effective description via scale decoupling. The parameter spaces suggest that the vacuum energy of active neutrino fields can naturally meet the observation of dark energy density, and a neutrino with a typical mass $\sim$ 10 meV $(10^{-3}$ eV) is expected. The normal ordering neutrinos are preferred by naturalness, and the neutrino mass window set by dark energy is 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV.
@@ I Introduction
Physics thrives on crisis Weinberg:1988cp , and the ultraviolet (UV) problem is one of the keys in the development of modern physics. In the context of the UV catastrophe in classical physics, the energy quanta were introduced by Planck Planck:1901tja while studying blackbody radiation problems, which led to the birth of quantum theory. In quantum field theory (QFT), the problem of UV divergences reappears, that is, loop corrections of a transition process are often UV divergent. It means that the transition amplitude obtained by Feynman rules (the physical input) is not directly equal to the physical result (the physical output). To extract the finite result from a UV-divergent input, the traditional approach is regularization & renormalization, that is, the UV divergence is firstly expressed by the regulator based on equivalent transformation and then removed by counterterms, with the mathematical structure of this route being $\infty$-$\infty$. This traditional approach is successful in dealing with logarithmic divergences in the standard model (SM). When we leave the logarithmic region and move on, there are UV problems for power-law divergences, i.e.,
@@
(b) The non-renormalizable Einstein gravity. When one tries to quantize Einstein gravity, an infinite number of counterterms are required for the graviton loops exhibiting power-law divergences.
@@
How do we deal with the above three UV problems? The traditional approach exhibits limitations in handling power-law divergences, that is, UV divergences under equivalent transformation demand mathematically precarious subtractions via counterterms, especially severe for power-law divergences. In this case, the free flow of ideas is important. Dirac DiracTheEO regarded renormalization as mathematically provisional in its handling of UV divergences; the physical laws chosen by nature were found to possess mathematical beauty. Feynman Feynman:1986er emphasized that mathematical self-consistency and rigor are essential in addressing UV divergences. In the vision of Dyson Dyson:1949ha and Schwinger Schwinger , future approaches to loop corrections would be based on finite quantities. Wilsonian effective theory Wilson:1971bg decouples scales via a physical cutoff that encodes UV effects into low-energy parameters. Hence, is it not time to explore a route that is intrinsically finite? For a new route, it first reproduces the renormalization results for logarithmic divergences, then addresses the three UV problems from power-law divergences.
For loops, assuming the physical contributions being local and insensitive to UV regions, a method of the UV-free scheme Jia:2023dub is introduced to derive loop contributions, which reproduces the renormalization results for logarithmic divergences. It takes a route of analytic continuation from the physical input $\mathcal{T}_{\mathrm{F}}$ to the physical output $\mathcal{T}_{\mathrm{P}}$ with a mathematical structure $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$,
@@
$\mathrm{Input}\bigg\{\!\!\begin{array}[]{c}\textit{Equivalent transformation}\,,\infty-\infty\\ \textit{Analytic continuation}\,,\,\,\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}\end{array}\!\!\bigg\}\mathrm{Output}\,$, (3)
@@
and UV divergences are systematically eliminated through $\mathcal{T}_{\mathrm{F}}\to\mathcal{T}_{\mathrm{P}}$ without counterterms. The loop contributions constructed through locally finite quantities are intrinsically free of UV divergences, as guaranteed by locality-based analytic continuation. For power-law divergences, this method inherently obviates UV subtractions and thus interprets the Higgs mass hierarchy problem within SM Jia:2023dub , and an application to Einstein gravity was investigated in Ref. Jia:2024lyy . Here we focus on the UV-divergent vacuum energy. The cosmological constant (dark energy) problem remains a fundamental challenge in physics Weinberg:1988cp ; SupernovaSearchTeam:1998fmf ; SupernovaCosmologyProject:1998vns ; Peebles:2002gy ; Padmanabhan:2002ji ; Copeland:2006wr ; Li:2011sd . In exploring the nature of dark energy, a central and unavoidable question is the vacuum energy density $\rho_{\mathrm{vac}}$ that corresponds to a cosmological constant ($\Lambda=8\pi G\rho_{\mathrm{vac}}$). Since the form of the vacuum energy ($\varpropto\!\int\!\!\frac{d^{3}k}{(2\pi)^{3}}\sqrt{k^{2}\!+\!m^{2}}$) resembles a loop integral, the UV-free scheme will be employed as a uniform treatment.
@@ II The vacuum energy in QFT
where $j_{i}$ is the spin of the field, and $g_{i}$ is the number of degrees of freedom. This is a quartic divergence, yet not a loop contribution. The UV-free scheme Jia:2023dub ; Jia:2024lyy , a physical analytic continuation of the UV-divergent integral, is adopted to evaluate the contribution of Eq. (4). The Feynman amplitude $\mathcal{T}_{\mathrm{F}}^{i}=\rho_{0}^{i}$ is the physical input, and the Feynman-like amplitude $\mathcal{T}_{\mathrm{F}}^{i}(\xi)$ here can be obtained with $\sqrt{k^{2}\!+\!m_{i}^{2}}$ replaced by $\sqrt{k^{2}\!+\!m_{i}^{2}\!+\!\xi}$. The physical transition amplitude $\mathcal{T}_{\mathrm{P}}^{i}$ is
@@
In the limit $m_{i}\to 0$, the result is zero for a massless field. [^1] For another type UV-divergence caused by the discrete sequence, i.e., the Casimir effect from bounded vacuum field modes, the Riemann zeta function $\zeta(-1)=-\frac{1}{12}$ can be adopted as the analytic continuation. The new energy scale $\mu_{\Lambda}$ is the scale characterizing vacuum energy, analogous to $\Lambda_{\mathrm{QCD}}$ for the strong interaction and $v$ for the electroweak sector.
Here is an interpretation of scales in QFT. In a scattering process, the full physical result (tree and loop levels) is independent of the free parameter $\mu$ used in perturbative expansions; it can be equivalently described by a tree-level process with an effective coupling at the scale $\mu_{\mathrm{eff}}$. The vacuum energy scale $\mu_{\Lambda}$ is analogous to this physical scale $\mu_{\mathrm{eff}}$. Unlike $\mu_{\mathrm{eff}}$, which varies with scattering energy, $\mu_{\Lambda}$ acts as a common reference scale for all field ground states, and may be fixed or slowly running with cosmic evolution.
@@ III The ripple description and neutrino field
The QFT vacuum, the ground state of quantum fields, contains zero-point fluctuations from the uncertainty principle. Their short-distance energy density is given by Eq. (6), showing that a field’s contribution to dark energy is not a UV problem. The central question thus becomes how mechanisms beyond current QFT suppress heavy-field contributions. The short-distance result cannot be directly identified with the macroscopic cosmological constant, implying a fundamental scale separation. In the Wilsonian framework Wilson:1971bg , this signifies a decoupling between microscopic zero-point energy and its macroscopic effect. We therefore adopt an effective description based on this scale decoupling, taking the total zero-point energy of free field modes as the physical baseline. Physical vacuum energy is then defined as the deviation from this baseline, caused by boundary conditions or interactions. We propose that dark energy originates from the cumulative effect of vacuum fluctuations mediated by carrier fields at the characteristic scale $\mu_{\Lambda}$. Here $\mu_{\Lambda}$ acts as the filtering scale for a field of mass $m_{i}$: fields with $m_{i}\lesssim\mu_{\Lambda}$ contribute actively, while those with $m_{i}\gg\mu_{\Lambda}$ remain inert. This picture is captured by the ripple description (Fig. 1): just as a net water displacement requires coherent ripples, a net physical contribution requires vacuum fluctuations to achieve macroscopic coherence at the scale set by $\mu_{\Lambda}$ via a fluctuation factor structuring the carrier field, which selects coherent contributions on macroscopic scales, analogous to how the Boltzmann factor selects energy quanta in thermal radiation. Given the Hubble parameter $H_{0}\approx$ 70 km s-1 Mpc-1 and dark energy at 71% of the critical density $3H_{0}^{2}/8\pi G$, the dark energy density is about $(2.3~\mathrm{meV})^{4}$, corresponding to $\mu_{\Lambda}\sim 10\!-\!100$ meV, indicating sensitivity to the neutrino field, in particular to neutrino mass. [^2] Other phenomenological proposals explore neutrino contributions via, e.g., the relic neutrinos with varying masses Fardon:2003eh ; Fardon:2005wc ; Takahashi:2005kw , the slow-roll quintessence with extended neutrino sector Hung:2000yg ; Barbieri:2005gj , the cosmological axion and neutrino deVega:2023fvw , etc. Considering Eq. (6) and dark energy density, a typical neutrino mass is $m_{\nu}^{\mathrm{tp}}\sim 2.3\times(32\pi^{2})^{\frac{1}{4}}\approx$ 9.7 meV (here $|\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}|^{\frac{1}{4}}\sim 1$). As a rough prediction, there should be a neutrino with a mass $\sim$ 10 meV.
@@
Now we turn to neutrino masses. The masses of three neutrinos are $m_{1}$, $m_{2}$ and $m_{3}$, and the mass differences between them can be extracted by neutrino oscillations. The mass-splitting results ParticleDataGroup:2024cfk are $\Delta m_{21}^{2}\approx 7.41\times 10^{-5}\mathrm{eV}^{2}$, $\Delta m_{31}^{2}\sim\Delta m_{32}^{2}\approx 2.437\times 10^{-3}\mathrm{eV}^{2}$ (normal ordering) and $\Delta m_{31}^{2}\sim\Delta m_{32}^{2}\approx-2.498\times 10^{-3}\mathrm{eV}^{2}$ (inverted ordering). Considering the naturalness with an effective mass $m_{\nu}^{\mathrm{tp}}$, the quasidegenerate spectrum with $m_{1}\simeq m_{2}\simeq m_{3}\gg\sqrt{|\Delta m_{32}^{2}|}$ is not favored. It is possible to give an estimate on the mass spectrum of neutrinos. Here we consider a specific implementation of the ripple description by a test fluctuation factor $e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ in Gaussian distribution (see the Appendix) to characterize the active contribution of a field, and $g_{i}$ together with this factor can be taken as an effective active degrees of freedom $g_{i}^{\ast}$ with $g_{i}^{\ast}=g_{i}e^{-m_{i}^{2}/\mu_{\Lambda}^{2}}$ ($\rho_{\mathrm{vac}}=\sum_{i}(-1)^{2j_{i}}g_{i}^{\ast}\frac{m_{i}^{4}}{64\pi^{2}}\log\frac{m_{i}^{2}}{\mu_{\Lambda}^{2}}$). For a field with a mass $m_{i}\gg\mu_{\Lambda}$, its contribution to the dark energy density is negligible due to the tiny value of $g_{i}^{\ast}$ (this is similar to the freezing of effective degrees of freedom of heavy particles at a low temperature in the early universe). The hierarchical spectra of neutrinos for normal ordering and inverted ordering with the required dark energy density are shown in Fig. 2 and Fig. 3, respectively. In Fig. 3, considering the naturalness of $\mu_{\Lambda}$, the inverted ordering is not favored. For the normal ordering shown in Fig. 2, there is a region that neutrino masses at given $\mu_{\Lambda}$ can naturally produce the physical vacuum energy required by the dark energy density, i.e., the parameter spaces forming a mass-energy window. Likewise, the neutrino mass window set by dark energy is 6.3 meV $\lesssim m_{1}\lesssim$ 16.3 meV, 10.7 meV $\lesssim m_{2}\lesssim$ 18.4 meV, 50.5 meV $\lesssim m_{3}\lesssim$ 52.7 meV, and the total neutrino mass is 67.5 meV $\lesssim m_{1}+m_{2}+m_{3}\lesssim$ 87.4 meV. This mass window can be tested by the future experiments.
@@
Let us give a brief discussion about the Hubble tension. A slightly higher $H_{0}=(73.04\pm 1.04$) km s-1 Mpc-1 Riess:2016jrr ; Riess:2021jrx from the nearby universe and a slightly lower $H_{0}=(67.4\pm 0.5$) km s-1 Mpc-1 Planck:2018vyg from the early universe indicate a possible Hubble tension (the tension may be alleviated Freedman:2024eph ). For $H_{0}\approx$ 67.4 km s-1 Mpc-1 and $H_{0}\approx$ 73 km s-1 Mpc-1, the corresponding dark energy densities are about $(2.24~\mathrm{meV})^{4}$ and $(2.37~\mathrm{meV})^{4}$, respectively. If the scale $\mu_{\Lambda}$ has a slow-running behavior and traverses the upper part of the mass-energy window in Fig. 2, the Hubble tension can be explained by this slow-running behavior, i.e., a running cosmological constant. If $\mu_{\Lambda}$ is fixed, new explanations and mechanisms need to be taken into account.